B-Waves. Starting with Maxwell’s equations, derive the 3-D wave equation for magnetic fields.

Gausss law for electric fields Gausss law for magnetic fields: Faradays law: (11-31a) (11-31b) (11-31c) OE Amperes law (11-31d)
Answer #1

\nabla \times B = \mu_0 \varepsilon_0 \frac{\partial \vec{E}}{\partial t}

Taking curl on both sides, we get

\nabla \times (\nabla \times B) = \mu_0 \varepsilon_0 \frac{\partial (\nabla \times \vec{E})}{\partial t}

\nabla (\nabla \cdot B) - \nabla ^2 B =  \mu_0 \varepsilon_0 \frac{\partial (\nabla \times \vec{E})}{\partial t}

From equation (11.31-b), \nabla \cdot B = 0 and from equation (11.31-c), \nabla \times \vec{E} = -\frac{\partial \vec{B}}{\partial t}

So,

\nabla^2 \vec{B} = \mu_0 \varepsilon_0 \frac{\partial^2 \vec{B}}{\partial t^2 }